OpenStudy (anonymous):

find the average rate of change from x=3 to x=15 dor the function f(x)=0.01(2)^x and select the correct answer below A.0.08 B. 12 C. 27.3 D. 327. 68 I'm leaning towards c or D maybe D

OpenStudy (anonymous):

@Rushwr

OpenStudy (anonymous):

ok so D

Vocaloid (vocaloid):

use the formula rate of change = [f(15)-f(3)]/(15-3)

OpenStudy (anonymous):

or a

OpenStudy (bibby):

nope

Vocaloid (vocaloid):

like I said, use the formula.. f(15) = ?

OpenStudy (anonymous):

not a or d

OpenStudy (bibby):

can you not follow instructions?

OpenStudy (anonymous):

Ok I thought it was C

Vocaloid (vocaloid):

please stop guessing... f(15) = ?

OpenStudy (anonymous):

ok then rate of change = [f(15)-f(3)]/(15-3) x=3 to x=15 if you use the formula the change in both should be 12

Vocaloid (vocaloid):

nope. f(15) - f(3) = ?

OpenStudy (anonymous):

15-3=12 f12

Vocaloid (vocaloid):

nope, that's not how it works. f(x)=0.01(2)^x f(15) = ?

OpenStudy (anonymous):

15f

Vocaloid (vocaloid):

no, please pay attention. f(15) means we let x = 15, and find f(x) what does f(x) equal when x = 15?

Vocaloid (vocaloid):

remember, f(x)=0.01(2)^x

OpenStudy (anonymous):

3? I'm terrible at math

Vocaloid (vocaloid):

x = 15 0.01(2)^x = ?

Vocaloid (vocaloid):

use a calculator if you are unsure

OpenStudy (anonymous):

0.08 if i evaluate it with x as 3

Vocaloid (vocaloid):

no, I just said that x = 15

OpenStudy (anonymous):

ok so evaluate 0.01(2)^x = ? so x is 15?

Vocaloid (vocaloid):

yes, evaluate 0.01(2)^x when x = 15

OpenStudy (anonymous):

327.68

Vocaloid (vocaloid):

good so f(15) = 327.63 f(3) = 0.08 now we can plug everything into the formula rate of change = [f(15)-f(3)] / (15-3) = ?

OpenStudy (anonymous):

I just get f

OpenStudy (anonymous):

Evaluate f?

Vocaloid (vocaloid):

f just means a function... we just found: f(15) = 327.63 f(3) = 0.08 what is f(15) - f(3)?

OpenStudy (anonymous):

12f

Vocaloid (vocaloid):

no...

Vocaloid (vocaloid):

f(15) is equal to 327.63 f(3) is equal to 0.08 what is f(15) - f(3)? the answer is NOT 12 f

OpenStudy (anonymous):

oh yeah ok

OpenStudy (anonymous):

I get 327.55

Vocaloid (vocaloid):

good, now divide that by (15-3)

OpenStudy (anonymous):

27.2958 repeating 3

OpenStudy (anonymous):

So I can just make it 30?

Vocaloid (vocaloid):

we can round that to 27.9

Vocaloid (vocaloid):

*27.3

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

because the 3 was repeating

Vocaloid (vocaloid):

and that's our answer, 27.3

OpenStudy (anonymous):

So C?

OpenStudy (anonymous):

thanks so much